OpenStudy (anonymous):

The area of a desk top is 8 ¾ square feet. If the length is 3 ½ft., find the width. 2 1/2 ft. 24 3/8 ft. 30 5/8 ft.

OpenStudy (anonymous):

Area = length x width so solve this equation: \[\huge 8\frac{3}{4}=(3\frac{1}{2})\cdot (width) \]

OpenStudy (anonymous):

solve for the width...

OpenStudy (anonymous):

mkay

OpenStudy (anonymous):

For me it's easier to simplify the values so they don't have fractions for example\[3\frac{1}{2}=3.5\] and \[8\frac{3}{4}=8.75\]

OpenStudy (anonymous):

@nicolelester123 , do you know what to do?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

do you know how to solve equations with one unknown for example: 3=2*M Would you know how to solve for M in the equation I posted?

OpenStudy (anonymous):

solve the equation for width....

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

It's the same thing here just solve for Width

OpenStudy (anonymous):

\[8.75=3.5*W\] Solve for W W=Width

OpenStudy (anonymous):

You seriously need to go back and read the book. If you don't know how to solve for an unknown you will have a lot of trouble in the future.

OpenStudy (anonymous):

ur answer is 2 1/2

OpenStudy (anonymous):

thank you(:

OpenStudy (anonymous):

ya im smart nah jk

OpenStudy (anonymous):

@colobikelawrider , please read the code of conduct for OS... don't just give answers. at least have some kind of justification...

OpenStudy (anonymous):

yeah you are! i'm stupid

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

lol... keep doing that and see what happens....

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Then whats the point of this site . When theres yahoo answers they answer faster and theres no rule of conduct